OpenStudy (anonymous):

What is the equation of the line passing through (8, –6) and perpendicular to the line y = x – 2 in slope-intercept form? Can someone please walk me through it? I really need to understand how to do this.

OpenStudy (anonymous):

I know how to write an equation of a line, but how do I make it perpendicular to another line?

OpenStudy (anonymous):

Well, if it's perpendicular (and parallel) are ALL ABOUT SLOPES :) Parallel lines have the same slope, while perpendicular lines have slopes which are negative reciprocals of each other :) What's the slope of y = x - 2 ?

OpenStudy (anonymous):

1 right?

OpenStudy (anonymous):

Right :) so the slope of the line you want to construct must be the negative of the reciprocal of 1, which is...?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

but why would it be changed to the negative reciprocal?

OpenStudy (anonymous):

I'm not really sure how to answer that :( I guess it just is :)

OpenStudy (anonymous):

Lol so if I have a slope of one and I am trying to do this type of question it is just the negative reciprocal?

OpenStudy (anonymous):

Well, it's the negative reciprocal because this new line you want is perpendicular to the given line :)

OpenStudy (anonymous):

Okay! Thanks! So what's next

OpenStudy (anonymous):

Well, you now have your slope, and a point. You know how to construct the line, given a point and a slope?

OpenStudy (anonymous):

Wait so my slope is -1 correct? And which point are we talking about now? The y-int or (8, -6) ?

OpenStudy (anonymous):

The point (8,-6), of course :)

OpenStudy (anonymous):

So then the y intercept would be positive 2 right?

OpenStudy (anonymous):

Hmm... given the slope m and a point (a ,b) on the line, the equation of the line is given by y - b = m(x - a) You have m = -1 (a,b) = (8, -6) Just plug in :)

OpenStudy (anonymous):

Ohhh okay! thank you so much!

OpenStudy (anonymous):

No problem :)